### v1

# Conceiving Naturally After IVF:

[…]
Conceiving Naturally After IVF:
the effect of assisted reproduction on obstetric interventions and child health at birth.
Fabio I. Martinenghi, Xian Zhang,
Luk Rombauts and Georgina M. Chambers
Introduction
It is estimated that 439,039 babies were born via assisted reproduction technology (ART) in 2018 alone (Chambers et al., 2021b). That same year, almost 5% of all the children born in Australia were conceived via in-vitro fertilisation (IVF, a form of ART; Newman et al., 2020), while estimates for the US and Europe are at 1.6% and 4.5%, respectively, and growing (De Geyter et al., 2018; Sunderam et al., 2018). In the context of this significant contribution of ART to global fertility, pregnancies conceived via ART are currently considered high-risk (see Berntsen et al., 2019, for a recent review of the literature). However, the causal evidence supporting this claim remains scant.
Indeed, while the characteristics of couples who use ART strongly correlate with poorer outcomes at birth111 This is not only true for multiple pregnancies, but also for singleton pregnancies. Singleton pregnancies are pregnancies where the mother is expecting only one child. Multiple pregnancies, such as twins and triplets, increase risks of adverse outcomes at birth for the mother and the children. They are common among IVF pregnancies although largely preventable by transferring a single embryo per each cycle. The Reproductive Technology Accreditation Committee of the Fertility Society of Australia and New Zealand requires to “minimise the incidence of multiple pregnancies”, which includes recommending a single embryo transfer per cycle, ensuring that no more than two embryos are transferred per cycle, and informing the patient of the risks associated with double transfers (Reproductive Technology Accreditation Committee, 2021). , both for mother (e.g. caesarean sections) and the baby (e.g. pre-term births), it remains unclear whether ART treatment independently increases the risk of such adverse outcomes. These characteristics include, for instance, advanced maternal age (Attali and Yogev, 2021; Lean et al., 2017; Berntsen et al., 2019), certain causes of infertility and comorbidities including obesity (Pasquali et al., 2007; Maheshwari et al., 2007). In relation to obstetric interventions, one concern is that the considerable emotional and financial investment implied by an ART pregnancy might exacerbate defensive medical behaviour, where physicians intervene more than necessary driven by the fear of medical liability (for evidence of its impact on c-section rates, see Dubay et al., 1999; Jena et al., 2015; Betrán et al., 2018; Bertoli and Grembi, 2019; Card et al., 2019; Laurita Longo et al., 2020).
In this paper, we study the effect of ART treatment on obstetric and perinatal outcomes. We measure its impact on the probability of pre-term birth and, more broadly, on the health of the baby. This includes the impact of ART on the baby’s gestational age at birth, weight at birth and its APGAR 1 and 5 scores (a measure of infant health). Finally, we study its impact on the rates of spontaneous labour (rather than induced), preterm spontaneous labour and c-section, which helps us understand whether ART pregnancies experience higher rates of interventions, holding all else constant.
Our approach leverages the stochastic nature of conception, including ART conception. Indeed, while factors such as age and reproductive health influence the probability of conception, chance (idiosyncratic factors) plays an important role. Fertile couples in their 20s have a $25\%$ chance of conceiving spontaneously in each menstrual cycle (ACOG, 2023), and a $41-47\%$ probability of conceiving via IVF per IVF cycle222 Estimates computed on 7 February 2024 via the online government-funded IVF prediction tools: YourIVFSuccess.com.au, based on Australian data, https://www.cdc.gov/art/ivf-success-estimator/index.html, based on US data, and https://w3.abdn.ac.uk/clsm/opis/tool/ivf1, based on UK data. The parameters set for the first prediction tool were: “Age of person intending to carry the pregnancy: 22”, “Sperm provider’s age: 22”, “Previous pregnancy: No”, “Main infertility diagnosis: I have not had any tests carried out, or I do not have medical infertility”, “Previous IVF treatment: No”. The outcome was ”having a baby in the 1st complete egg retrieval cycle” For the second tool the parameters were: “Age: 22 years”, “Weight: 110 lbs”, “Height: 5 feet”, “Number of IVF cycles used: I’ve never used IVF”, “Number of prior pregnancies: None ”, “Number of prior births: None”, “Egg source: My own eggs”, “Diagnosis: I don’t know/no reason”. The outcome was “cumulative chance of live birth after 1 retrieval and all transfers within 12 months.” For the third tool, the parameters were: “What is your age? 22”, “How many years have you been trying to conceive? 0”, “Have you been pregnant before? No”, “Do you have a problem with your tubes? No”, “Do you have an ovulation problem? No”, “Do you have a male factor fertility problem? No”, “Do you have an unexplained fertility problem? No”, “Do you plan to have IVF or ICSI? IVF” The outcome was “chance of having your first baby after 1 complete cycle of treatment.” . In either case, the residual variation in conception remains large and is plausibly idiosyncratic. The persistently low precision of clinical prediction models for IVF pregnancies, even the most recent ones (C-indices between 0.6-0.7 McLernon and Bhattacharya, 2023), is also consistent with some variation in IVF success being idiosyncratic.
Our strategy to isolate this variation has two components. First, we focus on births from women who have undergone ART treatment in Australia and subsequently given birth. We assign births whose conception depended on ART treatment to the treatment group, and assign births that were naturally conceived soon after ART treatment to the control group. Specifically, our treatment group includes singleton births from mothers who had a successful ART cycle, while our control group includes singleton births from mothers who had an unsuccessful ART cycle and conceived spontaneously between 3 and 12 months after the cycle (see Figure 3). This removes the selection bias associated with comparing the outcomes of fertile and subfertile women—populations that are different over both observable and unobservable characteristics.
Second, we flexibly adjust for key confounding variables, using both standard linear methods and flexible machine-learning-based methods (see Figure 2). Our linked administrative datasets allow us to observe both cycle-level information and some clinical history of the mother, which we use to control for key confounders such as age, parity, number of prior ART cycles, socio-economic characteristics and co-morbidities. Finally, we perform a number of sensitivity analyses, studying how unobserved confounders might affect our estimates (via Cinelli and Hazlett, 2019) and exploring how our results change using alternative sample restrictions and model specifications.
We do not find evidence that ART increases the risk of preterm birth or other adverse obstetric outcomes. In fact, the effects of ART on gestational age at birth, birth weight and APGAR 1 and 5 scores are not statistically different from zero at conventional levels. Our estimated null effects are precise, so that it is unlikely that clinically significant effect sizes exist and have not been detected 333 We can exclude the following (clinically small) effect sizes for the effect of ART on obstetric outcomes: effects between [-0.16, 0.10] weeks for gestational age at birth; between [-1.1, 1.9] p.p. for the risk of preterm birth; between [-0.07, 0.09] points for the APGAR1 score; between [-0.02, 0.1] points for the APGAR5 score; between [-14, 60] grams for the effect of ART on birth weight. . Moreover, we estimate that ART has a (precise) null effect on the risk of preterm spontaneous birth, an important adverse outcome that is not subject to medical intervention and indicates poor infant health. We also find that ART slightly decreases the risk of obstetric interventions, lowering the risk of a caesarean section (-5.1 p.p.) and increasing the rate of spontaneous labour (+4 p.p.).
We contribute to the literature on ART by proposing a new strategy that uses exogenous residual variation in ART success to identify the causal effect of ART. We focus on obstetric outcomes, but our strategy can be applied to other child- and mother-related outcomes. As Berntsen et al. (2019) point out, identifying the effect of ART on perinatal outcomes as separate from that of infertility remains a major challenge. Studies that use all spontaneously conceived births [^4] These are births from both fertile mothers and mothers with a history of infertility treatments. to construct a comparison group for ART birth (e.g. Farley et al., 2021) can only identify this effect by relying heavily on their observed controls, as the mothers of spontaneously conceived babies can be very different from ART mothers in terms of both observable and unobservable characteristics. Sibling studies such as Westvik-Johari et al. (2021)—where some siblings were conceived spontaneously, while others via ART—help controlling for maternal factors that do not change over time, but are confounded by maternal age, birth order and all the time-varying maternal factors. First, our approach implies conditioning on mothers with a history of ART. This eliminates an important source of selection bias, which is prevalent in studies that include births from fertile mothers. Second, it allows us to compare—to ART births—spontaneous births to similar mothers who also had recently undergone ART. This helps us build a credible counterfactual to conceiving via ART treatment while leveraging DML to flexibly control for a number of key confounders, including age and parity. Our strategy can be used in future research to study the effect of ART on other outcomes of interest.
The rest of the paper is structured as follows. Section 2 reviews the literature on ART and obstetric outcomes. Section 3 describes our sample and data sources, while Section 4 details the methods we use to estimate the effect of ART on obstetric outcomes. In Section 5, we conduct a descriptive analysis. In Section, 6 we report and discuss our findings. our sample. Section 7 closes with a review of the results.
Related literature
In this section, we start by defining assisted reproduction technologies and discussing related medical literature. We then discuss Health Economics research on ART.
Background and medical literature
Assisted reproductive technology (ART) is a mainstream treatment for infertility, with an estimated over two million cycles performed each year globally (Chambers et al., 2021c). There are over 500,000 ART-conceived births each year globally, accounting for more than 5% of total births in a number of high-income countries (European IVF Monitoring Consortium (EIM) et al., 2022; Choi et al., 2022). ART treatment, otherwise known as IVF, involves a sequential procedure where (i) the oocyte (egg) and sperm are joined together outside the body in a specialised laboratory, (ii) the fertilised egg (embryo) is allowed to grow in a protected environment for 4-6 days, and (iii) the embryo is transferred into the woman’s uterus with the hope of a pregnancy. Perinatal outcomes are reported to be poorer in ART-conceived pregnancies compared to pregnancies conceived after a spontaneous conception, and this might be attributed either to the underlying infertility, to the ART treatment, or both (Berntsen et al., 2019). Several meta-analyses and cohort studies have compared perinatal outcomes for ART-conceived singletons with spontaneously-conceived singletons and found that ART-conceived singletons were at risk of poorer perinatal outcomes, including a higher risk preterm birth (Helmerhorst et al., 2004; Jackson et al., 2004; Pandey et al., 2012; Marino et al., 2014; Qin et al., 2017), low birth weight (Helmerhorst et al., 2004; Jackson et al., 2004; Pandey et al., 2012; Marino et al., 2014; Qin et al., 2017), and perinatal mortality (Helmerhorst et al., 2004; Jackson et al., 2004; Pandey et al., 2012; Marino et al., 2014; Qin et al., 2017; Farley et al., 2021). Moreover, previous studies reported that compared to spontaneously-conceived pregnancies, ART-conceived pregnancies are associated with increased obstetric complications, such as caesarean section (Helmerhorst et al., 2004; Pandey et al., 2012; Buckett et al., 2007), induction of labour (Pandey et al., 2012), gestational diabetes (Pandey et al., 2012), and gestational hypertension (Pandey et al., 2012; Opdahl et al., 2015; Thomopoulos et al., 2013).
However, given that these obstetric complications are relatively rare the absolute risk due to ART remains quite small (Pandey et al., 2012). A major challenge for research into the obstetric and perinatal outcomes following ART is how to separate the contribution of underlying infertility from the ART treatment per se. Currently, it remains unclear to what extent these adverse obstetric and perinatal outcomes can be attributed to the ART treatment, rather than the underlying infertility and other risk factors correlated with ART use. Our study addressed this issue.
Health Economics literature
There is growing interest in fertility treatments among economists, as the extension of the fertility window they bring has the potential to affect the childbirth and career decisions of women, among other things. In particular, US states have progressively mandated that private insurers offer coverage for fertility treatment. Such mandates have informed a growing body of literature about how the resulting increased access to fertility treatments (Hamilton and McManus, 2012) has increased the birth rates for women over age 35 (Schmidt, 2005, 2007), increased multiple-birth rates for all women (Zaresani and Schmidt, 2023), delayed the timing of marriage (Abramowitz, 2014) and, the extent to which existing disparities in treatment use have increased (Bitler and Schmidt, 2006; Buckles, 2013), with the largest utilization of in fertility treatments found in older, more educated women (Bitler and Schmidt, 2011), and the extent to which access to ART treatment have decreased the rate of child adoption
Hamilton et al. (2018) study these insurance mandates formally, constructing and estimating a dynamic model for IVF patients’ choices. To reduce patient welfare while lowering medical costs, they recommend policies that increase insurance coverage while charging more for multiple births (twins, triplets, etc.), as they have medical costs much higher than singleton births. More broadly, Bhalotra et al. (2022) provide evidence that, compared to singletons, twin births not only decrease maternal and child health, they also lead to steeper decreases in maternal earnings and lower subsequent fertility. Their study leverages a Swedish policy mandating that IVF embryo transfers must be single, rather than multiple, hence lowering the chances of a multiple birth for IVF prospective mothers.
An international study of 30 countries that systematically quantified the impact of consumer cost on assisted reproduction technology (ART) utilization and number of embryos transferred found that the relative cost that consumers pay for ART treatment (after taking account of insurance, wages and disposable income) predicts the level of access to treatment as well as the number of embryos transferred (Chambers et al., 2014). After controlling for other explanatory variables, affordability had a strong and robust association with utilization, with a 10-percentage-point decrease in affordability predicted to, on average, decrease ART utilization by 32%. Affordability was also independently associated with the number of embryos transferred with less affordability treatment leading to more risky embryo transfer practices.
Affordability and access have not been an issue for Israeli citizens since 1994, when the government made IVF free to all citizens up to the birth of two children Gershoni and Low (2021b, a). Comparing men’s outcomes to those of women, they find that this reform delayed the age at first marriage for women and their age at first birth. It also led to an increase in the number of women who have completed college and graduate education. This latter finding is consistent with Kroeger and La Mattina (2017), who find US evidence that increased coverage for fertility treatments makes college-educated women more likely to be in professional occupations and to complete a professional degree.
A different approach is taken by Lundborg et al. (2017), who do not study the impact of IVF directly, and instead use IVF success as an instrument for childbearing. They find that childbearing (i.e. a successful IVF cycle) leads to a persistent decrease in women’s earnings. They work less when the children are young and move to lower-paid jobs closer to home as the children get older—thus earning less.
Finally, our study contributes to the literature on “defensive medicine” and c-sections. This literature finds robust cross-country evidence that part of the variation in caesarean rates can be explained by the physician’s fear of litigation (Dubay et al., 1999; Jena et al., 2015; Betrán et al., 2018; Bertoli and Grembi, 2019; Card et al., 2019; Laurita Longo et al., 2020). Given the findings of this literature, one could expect this phenomenon to be even more pronounced in IVF pregnancies, which not only tend to be high risk, but might have also required a considerable financial investment. Our study addresses our question, studying whether ART leads to changes in caesarean rates.
Data
This study was approved by the NSW Population and Health Services Research Ethics Committee (2017/HRE1202), the ACT Health Human Research Ethics Committee (ETH.2.218.032), the Calvary Public Hospital Human Research and Ethics Committee (3-2018), and the Australian Institute of Health and Welfare Ethics Committee (AIHW) (EO2017/4/420),
Our cohort was extracted from a newly created medically assisted reproduction (MAR) data linkage, described in detail elsewhere (Chambers et al., 2021a; Choi et al., 2022). This linkage contains longitudinal demographic, treatment, and outcome data for women who underwent ART treatment in New South Wales (NSW) and Australian Capital Territory (ACT) between 2009 and 2017. Combined, NSW/ACT represents Australia’s most populous jurisdictions with more than 8.5 million residents and nearly 100,000 births per year (Australian Bureau of Statistics 2021).
Briefly, the MAR data linkage includes all births (mother and baby records) recorded in NSW and ACT perinatal data collections (PDCs) that are linked to the Australian and New Zealand Assisted Reproduction Database (ANZARD), a clinical quality registry of all ART treatments and outcomes performed across Australia and New Zealand. Given the licencing requirement for fertility clinics to submit data to ANZARD, complete ascertainment can be assumed. Over 96% of ANZARD records were able to be linked to jurisdictional datasets and over 94% of births recorded in ANZARD also had a birth recorded in the PDCs with very high concordance between birth outcomes (Chambers et al., 2021a). In addition, the linkage includes data from the NSW and ACT Admitted Patient Data Collection (APDC) and the Registry of Births, Deaths and Marriages (RBDM).
Study cohort
From the wider MAR data linkage, we make use of the ANZARD and PCD datasets, which give us detailed information on ART treatment history of mothers and on their babies’ perinatal outcomes.
Our sample covers all the births that occurred between 2009-2017 in New South Wales. We drop births in the Australian Capital Territory as they are only available for public hospitals—contrary to the NSW data, which also covers private-hospital births. We also drop mothers with tubal disease, as they would not be able to conceive without ART and thus would not be eligible for the control group.
Treatment and control groups
We construct the treatment group by selecting all ART-conceived births resulting from ART treatment received in 2009-2017. We construct the control group by selecting all births conceived spontaneously (i.e. ART-independent births) within a 3-to-12-month window after a failed ART cycle received in 2009-2017. The 3-to-12 month windows was chosen to make the treatment and control groups as similar as possible over unobserved characteristics and to minimise the risk of misclassification between ART and spontaneous births. The upper bound of the window needs to strike a balance between bias and variance. The smaller the window, the greater the similarity between treatment and control groups, at the price of fewer observations. This initial sample includes 21,102 births, of which 19,646 are in treatment group and 1,456 are in the control group.
Then, we drop births with at least one missing observation, 7477 treatment births (38.1%) and 521 control births (35.8%) leaving us with 12,169 treatment births and 935 control births555 These values change slightly across our outcomes at the estimation stage as, when we run the analysis for each outcome, we drop the births for which that outcome is missing.. These missing rates are mostly driven by the gestational diabetes and diabetes mellitus variables, of which 32% of the observations are missing. These are important variables, either confounders or closing a back-door path, and excluding them would likely bias our estimates. Our other covariates have less than 0.5% of missing observations with the exception of “smoking”, which has 3.6%.
The key issue informing the width of the conception window is that prospective mothers experiencing an unsuccessful ART cycle tend to schedule the next one at a close date, if they decide to schedule one666 In our sample of subfertile and ART-conceived singletons, the median time between unsuccessful cycles is 97 days. The 25th percentile is 56 days and the 75th percentile is 322 days. . Hence, if we compare a mother that has a successful ART cycle with one that conceived spontaneously a long time after a failed cycle, we are comparing a woman that might have undertaken several more cycles, had she been unsuccessful, with a woman that did not do that. These two women are likely different in important ways, and this would likely be a source of selection bias in our analysis. Figure 3 can help visualising how the control group is constructed.
Outcomes and covariates
Our outcomes and covariates of interest are presented in Table LABEL:tab:des. These covariates have been selected as they are key confounders777 While pre-eclampsia, gestational hypertension and gestational diabetes are not confounders, it is important to include them as they might be blocking a backdoor path or mediate the effect of ART on obstetric outcomes, the latter meaning that they can both have an impact on our outcomes and on the probability of a successful ART cycle—the treatment.
We complement the information in Table LABEL:tab:des with a few remarks. First, APGAR scores are standard measures of newborn health at birth and, as such, are strong predictors of neonatal death and neurological disorders (e.g. Persson et al., 2018; Cnattingius et al., 2020), among other adverse outcomes. Lower values imply lower newborn health.
Second, the confounder recording the number of previous failed ART cycles was based on the number of autologous ART cycles888 An autologous ART cycle is an ART cycle where the woman uses her oocyte or embryo, rather than one donated by a third party. In other words, we excluded donor cycles. , fresh or thaw—excluding freeze-all oocyte and/or embryo cycles 999 These are cycles where a woman freezes her oocytes or embryos for future use. . Women donating or freezing their oocytes and/or eggs were excluded because they weren’t at risk of pregnancy.
Third, the Socio-economic Indexes for Areas (SEIFA; ABS, 2011) variable used is the Index of Relative Socio-economic Disadvantage (IRSD). It is constructed by the Australian Bureau of Statistics and calculated using 2011 Census data, at the base level of Statistical Area Level 1 (SA1). A number of variables are used to measure determinants of disadvantage at the SA1 level, such as the proportion of unemployed or of single parents. Then principal component analysis is used to determine the variable weights in a data-driven way, resulting in the index. Finally, the index is divided into quintiles, which we use in our analysis. We are able to assign this variable thanks to SA1 information provided in the Perinatal Data Collection dataset.
Table 1: Variable description and source
Variable Source Description
Outcomes
Gestational age at birth PDC No. of completed weeks of gestation at the time of birth
Preterm birth PDC Equal to one if the baby is born after $<37$ weeks of gestation and zero otherwise
Spontaneous labour PDC Equal to one if labour was not induced (spontaneous) and zero otherwise
Preterm Spont. labour PDC Equal to one if the labour was spontaneous and the birth preterm, and zero otherwise
C-section PDC Equal to one if the baby was delivered via c-section and zero otherwise (i.e. vaginal delivery)
APGAR 1 & APGAR 5 PDC Count variables with values between 0-10, measuring newborn health at 1 and 5 minute from birth, respectively.
Birth weight PDC Weight at birth, in grams
Covariates
Age PDC Mother’s age at time of birth
Parity PDC Mother’s number of previous pregnancies greater than 20 weeks gestation
No. of failed ART cycles ANZARD No. of past failed ART cycles
Maternal YOB PDC Maternal year of birth
SEIFA index ABS Quintiles of the 2011 Census Index of Relative Socio-economic Disadvantage (ABS, 2011)
Remoteness of area PDC 4 levels of remoteness of mother’s dwelling, each encoded into one binary variable. Derived from the Accessibility/Remoteness Index of Australia (ARIA+) (ABS, 2001)
Born in Australia PDC Binary variable equal to one if the mother was born in Australia and zero otherwise
Diabetes mellitus PDC Binary variable for maternal diabetes mellitus
Chronic hypertension PDC Binary variable for maternal chronic hypertension
Pre-eclampsia PDC Binary variable for maternal proteinuric gestational hypertension
Gestational hypertension PDC Binary variable for maternal non-proteinuric gestational hypertension
Smoke PDC Binary variable equal to one if the mother smoked at any time during the pregnancy
Endometriosis ANZARD Binary variable equal to one if the treating clinician thinks subfertility is due endometriosis
Male infertility ANZARD Binary variable equal to one if the treating clinician thinks subfertility is due to a male factor problem
Other infertility ANZARD Binary variable equal to one if the treating clinician thinks subfertility is due to other female factors apart from tubal disease and endometriosis
Unexplained infertility ANZARD Binary variable equal to one if the treating clinician thinks there is clinical subfertility without apparent explanation
Methods
In this section, we define the ideal experiment that our observational analysis approximates (Section 4.1), followed by a presentation of the empirical methods we use (Section 4.2).
Ideal experiment
Our study can be conceptualised as approximating a two-stage randomised experiment. As shown in Figure 1, in the ideal experiment, we would take the whole population of prospective mothers undergoing an ART cycle and determine the success of their cycle at random, thus dividing them into two groups. Then, we would randomly “assign” a spontaneous conception to each of these two groups, leaving us with four groups (A, B, C, and D in Figure 1). From these, we exclude Group C as a spontaneous conception shortly after a successful ART cycle is almost impossible and we do not observe it. By this we mean that within our 3-to-12-month window from a successful ART cycle—hence resulting in a live birth 9 months after conception—there are only 3 months for an ART mother to conceive again and spontaneously. We also exclude Group B as it does not lead to a conception—our obstetric outcomes are not defined for this group. Instead, we focus on Groups A and D. Notice that here there are two “treatments”: ART success and spontaneous conception after a failed ART cycle.
If this this two-step randomisation model were the correct model of the world, we should observe excellent balance between our groups of interest—Group A (spontaneous conception) and Group D (ART conception). We do not impose this strong assumption though. Instead, we make a conditional exogeneity/independence claim. This means assuming that the treatment assignments are independent of our outcomes of interest after controlling for the relevant confounders—variables causing changes in both treatment and outcome.
Given that our dataset includes key confounders and covariates, we see this conditional independence assumption as credible. Nonetheless, as no study can credibly claim to control for all confounders, we study how violations of the conditional independence assumption affect our estimates in Section 6.3.
Figure 1: Diagram of the ideal experiment
Empirical specifications
Once we have defined our cohort, we take two approaches to estimate the effect of ART on obstetric outcomes. First, we estimate the following model via OLS:
$Y_{i}=\tau{D}_{i}+\mathbf{X}_{i}^{\prime}\bm{\beta}+\varepsilon_{i},$ (1)
where, $Y_{i}$ is the outcome variable, $D_{i}$ is the treatment variable, and matrix
$\mathbf{X}_{i}=\left(X_{1i},\ldots,X_{pi}\right)$
consists of the covariates listed in Table LABEL:tab:des—which enter the equation linearly—and $\varepsilon_{i}$ is the error term. Assuming that D is conditionally exogenous/independent, i.e.
$\mathbb{E}\left[\varepsilon_{i}\mid D_{i},\mathbf{X}_{i}\right]=0,$
and that $\tau$ is homogeneous across the treatment and control groups, then $\tau$ measures the average treatment effect (ATE) of treatment $D_{i}$ on outcome $Y_{i}$.
Second, we take a double machine learning approach and estimate the treatment effects using an Interactive Regression Model (Chernozhukov et al., 2018). Given a binary treatment variable, the Interactive Regression Model estimates fully heterogeneous average treatment effects. Here, “fully heterogeneous” means that each of the two response curves —the potential outcomes as a function of $\mathbf{X}$—is allowed to be a different nonparametric function. Vectors $\mathbf{Y_{i}}$ and $\mathbf{D_{i}}$ are modelled such that
${Y_{i}}=g_{0}({D_{i}},\mathbf{X_{i}})+{U_{i}},$ $\mathbb{E}({U_{i}}\mid\mathbf{X_{i}},{D_{i}})={0},$ (2)
${D_{i}}=m_{0}(\mathbf{X_{i}})+{V_{i}},$ $\mathbb{E}({V_{i}}\mid\mathbf{X_{i}})={0}.$ (3)
We still target and estimate the ATE, here written as:
$\tau=\mathbb{E}\left[g_{0}({1},\mathbf{X}_{i})-g_{0}({0},\mathbf{X}_{i})\right].$
As mentioned above, in this framework we do not need to assume homogeneous treatment effects. Variables $\mathbf{X}_{i}$ are allowed to affect both the outcome $\mathbf{Y}$ and the treatment $\mathbf{D}$, each via a different function—$m_{0}(\mathbf{X})$ and $g_{0}(\mathbf{X})$, respectively. In particular, $m_{0}(\mathbf{X})$ is the propensity score and as such, its estimate can be inspected and used to test assumptions such as common support. This can help dissipate the black-box-type concerns typical of machine-learning applications.
Moreover, the Interactive Regression Model identifies the ATE under the unconfoundedness (or conditional independence) assumption (Rosenbaum and Rubin, 1983). This assumption has three parts. First, it assumes that the probability of a birth being conceived via ART (vs. spontaneously) does not depend of the probability of other babies being conceived via ART (no network/spillover effects). Second, it assumes that the probability of a birth being conceived via ART does not depend on their potential outcome (no selection into treatment). This means that a birth’s probability of being conceived via ART cannot depend, for instance, on how likely the birth is born pre-term. Third, it assumes that no birth is conceived via ART with certainty (common support assumption).
Selection into treatment is generally a major concern in studies on ART, as women are not randomly allocated to ART treatment, but rather choose to undertake ART treatment based on their fertility status. We deal with this issue by only selecting mothers that have undergone at least one ART cycle. Hence, this assumption is not of concern in our study. The same argument applies to point the no network/spillover effects assumption. Finally, the partially stochastic nature of ART success (i.e. it is not a deterministic process), combined with an inspection of Figure 5, should convince that the common support assumption holds.
Figure 2: Diagram summarising the study
Figure 3: Diagram illustrating the control-group construction
Our final step involves a formal discussion of omitted variable bias. Using the tools developed in Cinelli and Hazlett (2019), we examine how sensitive our OLS estimates are to the omission of confounding variables—typically because they are not observed by the analyst.
Descriptive analysis
We start from a descriptive analysis of our sample. Table LABEL:tab:tab:cov_balance_anzard_12 reports difference in means across treatment groups. We find a clinically important difference in maternal age, which is higher by 1.9 years (36.6-34.7 years) in mothers that spontaneously conceived after a failed ART —at an age where maternal fertility is in significant decline. This is also reflected in a lower maternal year of birth. In our sample, mothers who spontaneously conceived after ART are less likely to be at their first cycle ($0.25-0.36=-0.10$ p.p.) and have had more unsuccessful cycles before their pregnancy ($3.1-1.6=1.5$ cycles). A last noticeable difference is in the prevalence of male infertility, which is 7 p.p. higher in mothers of ART-conceived babies. The two groups are clinically similar in other respects, such as number of past pregnancies, socio-economic disadvantage, co-morbidities and cause of infertility.
Figure 4 plots the densities of a selection of covariates. This is to reassure us that the distribution of the covariates are similar across the two treatment groups, so that the observed differences in means are simply capturing a distribution shift.
Figure 4: Densities for non-binary covariates
Results
In this section, we take our models to the data and estimate the effect of ART on obstetric outcomes. In Section 6.1, we estimate a linear regression model via OLS. This is best approach under the assumption that the model is approximately linear, and sometimes even beyond linearity (Hansen, 2022). Because we do not know if the true data-generating process is approximately linear or not, in Section 6.2 we estimate our model flexibly using double machine learning. Finally, Section 6.3 studies how omitted variable bias (and misspecification) in our linear regression model could affect our results.
Average treatment effects for binary outcomes (see Table LABEL:tab:des) should be interpreted as percentage point changes. For instance, a coefficient of 0.046 for spontaneous labour means that we estimate that ART treatment decreases the probability of a spontaneous labour by 4.6 percentage points (p.p.). The coefficients for gestational age, APGAR 1 and 5, and birth weight should be interpreted as shifts in levels, i.e. a 0.05 coefficient means that ART treatment increases gestational age/APGAR1/APGAR5/birth weight by 0.05 weeks of gestation/APGAR1 points/APGAR5 points/grams. The above applies to both OLS and DML estimates.
Linear regression
Table LABEL:tab:tab:ols reports the OLS estimates of Equation 1 under our full sample (column 1), and under further sample restrictions (columns 2-6)  [^10] When restricting the sample, we adjust the model accordingly, so that, for instance, we drop the Number of previous pregnancies variable if we are conditioning on having no previous pregnancies.. In column 2, we only include births whose mother had no previous pregnancies longer than 20 weeks of gestation (nulliparous) nor previous unsuccessful ART cycles. In column 3, we only include births whose mother had no previous pregnancies longer than 20 weeks of gestation. In column 4, we only include, in the control group, births whose mother had a single previous unsuccessful ART cycles, and in the treatment group, ART births conceived at the first ART cycle. Finally, in column 5 we drop all our covariates and instead stratify by our most important covariates. More precisely, we estimate a difference in means across treatment groups in a sample restricted to nulliparous mothers at their first ART cycle, aged less than 40 and without comorbidities (selected primips).
The estimated treatment effects are, in general, clinically small [TO BE CONFIRMED BY THE CLINICIAN]. Focusing on our baseline estimates (Column 1), we find that ART has no significant effect on the rate of preterm births, on the rate of spontaneous preterm births, and on the average gestational age at birth—neither at the statistical nor at the clinical level [TO BE CONFIRMED BY THE CLINICIAN]. Taking preterm birth as an example, the 0.008 coefficient, with a 0.01 standard error, indicates that our best estimate for the effect of ART treatment of preterm birth is 0.8 p.p., but we cannot exclude that the true estimate is equal to zero at the 95% confidence level. The implied confidence interval is [$-0.0116,0.0276$], so that we cannot exclude effects ranging from -1.16 p.p. to +2.76 p.p [^11] Confidence intervals, while omitted in Table LABEL:tab:tab:ols due to space constraints, are reported for the double machine learning estimates, which are very similar to the OLS estimates and are reported Tables LABEL:tab:tab:dml_forest_ate_anzard_12 and LABEL:tab:tab:dml_forest_ate_anzard_12_trim_15.
The same applies to birth weight and to the APGAR 1 and 5 scores, whose estimated effects are not statistically significant and are clinically small [TO BE CONFIRMED BY THE CLINICIAN]. To give an idea of the magnitudes of these statistically insignificant estimates, the effect of ART on APGAR1 is estimated to be $-0.013$ points, on a 1-to-5 scale, while the effect on gestational age at birth is estimated to be $-0.057$ weeks, which less than 10 hours. As shown by the p-values in the square brackets, these estimates are not statistically significant at the 1% level or below, despite small standard errors. Considering that gestational age is measured in weeks, the APGAR scores are integers between zero and five and that birth weight is measured grams, we can notice how imposing additional restrictions on the sample does not shift the estimates in a clinically significant way for the most part [TO BE CONFIRMED BY THE CLINICIAN]. Understandably, our estimates’ standard errors and statistical significance varies as the sample gets smaller.
Moreover, ART increases the chances of a spontaneous labour by 4.6 p.p., implying that ART mothers are less likely overall to be subject to a labour induction than control mothers, who conceived spontaneously after a failed ART cycle. This reduction in interventions in also found at the mode-of-delivery level, where ART reduces the overall chances of a C-section by 5 p.p. Similarly to the estimated impacts on gestational age, APGAR scores and birth weight, imposing further sample restrictions increases our standard errors but does not have a clinically significant impact on our estimated treatment effects.
The estimates from restricted samples used in Columns 2-5 are largely consistent with our baseline estimates in Column 1. The coefficients for c-section and spontaneous labour, which are statistically significant in Column 1, remain consistent in their sign (with one exception). They do lose statistical significance though, as the sample size drops—which is an issue particularly for the control group. The coefficients for statistically The coefficients for the other outcomes, not significant in Column 1, remain mostly small, while becoming more noisy due to reduced sample sizes.
Double machine learning
We start the analysis of our double machine learning model by inspecting its propensity score estimates, which the model uses to weight each observation and ultimately estimate the treatment effects. Figure 5 shows the estimated propensity score by treatment group. This is useful for two reasons.
Figure 5: Propensity score density by treatment group
First, we can check whether the common support assumption is met. This implies that no unit should have a score of one or zero, which we observe. Units close to zero or one would be of concern too as they would be assigned a high weight, and we have some observations between a cluster of observation around 0.95, with a maximum of 0.976. We take this into account in our analysis by trimming all observations outside the $[0.15,0.85]$ propensity score interval.
Second, Figure 5 allows us to inspect the degree of balance between the two treatment groups. We can see that the two distributions overlap over most of the support, with the notable exception of the $[0.85,0.976]$ interval. Again, we take this into account by trimming all observations outside the $[0.15,0.85]$ propensity score interval.
Table LABEL:tab:tab:dml_forest_ate_anzard_12 reports the baseline DML treatment effect estimates, while Table LABEL:tab:tab:dml_forest_ate_anzard_12_trim_15 reports the DML estimates after trimming the propensity score at 0.15 and 0.85. Not only the baseline DML estimates are robust to trimming, they are also virtually identical to the OLS estimates (Table LABEL:tab:tab:ols , Column 1), as can be visualised in Figures 7 and 6.
Figure 6: ATE estimates across outcomes and models
Figure 7: Birth weight ATE estimates across models
Omitted variable bias analysis
This similarity between the estimates from the linear regression and the flexible DML models is evidence that the data-generating process is not highly non-linear and, conversely, that it is approximately linear. Hence, the linear regression estimates should be taken seriously. One last concern remains, that some important confounders have not been included in our model and would drastically change our results if included.
We study this possibility graphically using the tools developed in Cinelli and Hazlett (2019). In Figure 9, we show what would happen in to our baseline linear regression estimate (left column) and t-static (right column) if we were not controlling for one or more confounders once/twice/thrice as strong as maternal age (top row), number of past pregnancies (middle row), or number of past cycles (bottom). While the estimated coefficient (t-value) is depicted as a black triangle, the hypothetical coefficients (t-values) under the x1, x2 and x3 scenarios are represented by red rhombi, progressively apart from the triangle.
In a coefficient plot, a rhombus crossing the red line (at zero) means that the coefficient would change sign under that scenario. We see this for birth weight (Figure 8), in presence of an unobserved confounder three times as strong as the parity variable. In this case though, the rhombus is almost on the zero line and hence treatment effect would be virtually equal to zero.
In a t-statistic plot, a rhombus crossing the red line (at 1.96) means that either that (i) a statistically insignificant coefficient (at the 5% level) would become significant in that scenario or (ii), vice versa, that a statistically insignificant coefficient would become significant. We do not see this case, as no t-statistic associated with a statistically insignificant baseline coefficient crosses the 1.96 line. We see case (ii)for spontaneous labour in presence of an unobserved confounder three times as strong as Maternal age, which would make the ATE coefficient insignificant while moving it closer to zero, at 3 p.p. The plots for the remaining outcomes are included in the appendix. The plots for the C-section outcome resemble those for Spontaneous labour, while the plots for the APGAR scores and for Gestational age at birth resemble those for Birth weight.
Overall, these plots do not change the interpretation of our results. Even after using the three strongest confounders as a benchmark for unobserved confounders, with still find evidence that the effect of ART on obstetric outcomes is either small (Spontaneous labour, C-section) or precisely zero (gestational age outcomes, APGAR score outcomes and Birth weight). These results, taken as a whole, are evidence that there no clinically significant differences between ART babies and spontaneously-conceived babies, at least in subfertile mothers.
Refer to captionATE coefficientPartial $R^{2}$ of confounder(s) with the outcomeMaternal age at birthRefer to captiont-statisticRefer to captionPartial $R^{2}$ of confounder(s) with the outcomeNo. of past pregnancies (parity)Refer to captionRefer to captionPartial $R^{2}$ of confounder(s) with the treatmentPartial $R^{2}$ of confounder(s) with the outcomeNo. of past failed cyclesRefer to captionPartial $R^{2}$ of confounder(s) with the treatment
Figure 8: Sensitivity of birth-weight ATE estimate to unobserved confounding
Refer to captioncoefficientPartial $R^{2}$ of confounder(s) with the outcomeMaternal age at birthRefer to captiont-valueRefer to captionPartial $R^{2}$ of confounder(s) with the outcomeNo. of past pregnanciesRefer to captionRefer to captionPartial $R^{2}$ of confounder(s) with the treatmentPartial $R^{2}$ of confounder(s) with the outcomeNo. of past failed cyclesRefer to captionPartial $R^{2}$ of confounder(s) with the treatment
Figure 9: Sensitivity of pre-term-birth ATE estimate to unobserved confounding
Refer to captioncoefficientPartial $R^{2}$ of confounder(s) with the outcomeMaternal age at birthRefer to captiont-valueRefer to captionPartial $R^{2}$ of confounder(s) with the outcomeNo. of past pregnanciesRefer to captionRefer to captionPartial $R^{2}$ of confounder(s) with the treatmentPartial $R^{2}$ of confounder(s) with the outcomeNo. of past failed cyclesRefer to captionPartial $R^{2}$ of confounder(s) with the treatment
Figure 10: Sensitivity of spontaneous-labour ATE estimate to unobserved confounding
Conclusion
We study the effect of ART on obstetric outcomes. We use the outcomes of babies conceived spontaneously shortly after a failed ART cycle to build the counterfactual outcome, i.e. the outcome if that same baby was instead conceived via ART. To do this, (i) we select babies born 3-12 months after a failed ART cycle as the control group and ART-conceived babies as the treatment group; (ii) we estimate the average treatment effect of ART on preterm birth, spontaneous preterm birth, gestational age at birth, APGAR score and birth weight, using a linear model and a more flexible double machine learning model; (iii) we study omitted variable bias formally, using a selection of strong confounders as a benchmark for unobserved confounders. To our knowledge this is the first study to evaluated obstetric outcomes by constructing cohorts from women who have undergone ART, but who have conceived spontaneous or as a result of treatment, thus removing much of the bias and confounding associated with previous studies.
We find that conceiving a baby via ART does not affect the risk of preterm birth, preterm spontaneous labour, the gestational age at birth, the birth weight of the baby nor their APGAR scores, in subfertile mothers. Our estimated null effects are precise, leaving us unable to rule out only ART effects that are clinically small 121212 We can exclude the following effect sizes for the effect of ART on obstretric outcomes: effects between [-0.16, 0.10] weeks for gestational age at birth; between [-1.1, 1.9] p.p. for the risk of preterm birth; between [-0.07, 0.09] points for the APGAR1 score; between [-0.02, 0.1] points for the APGAR5 score; between [-14, 60] grams for the effect of ART on birth weight. . Moreover, we find that ART has a small negative impact (${}-5p.p.$) on the chance of a c-section and of an induction of labour. While the sign of these effect is robust to large levels of unobserved confounding, such large levels of confounding effects would further reduce the magnitude of these estimates and make them statistically insignificant.
Overall, these results support the hypothesis that ART does not independently affect obstetric risk in subfertile women. This study indicates that while ART treatment is an indicator of poorer obstetric and perinatal outcomes in singleton pregnancies (i.e. because of coexisting risk factors such as advanced maternal age and conditions associated with infertility), ART treatment per se does not appear to be an independent risk factor. These findings should be reassuring for women undergoing ART and fertility clinicians and hopefully provide context around ART-conceived pregnancies.